# Immunization against harmful fine-tuning attacks

### v1

@@ Appendix A Analysis of Misaligned Models in the wild
This section shows multiple cases where openly available LLMs have been intentionally trained for harmful purposes that could be considered harmful, categorises types and methods of these attacks and showcases that control techniques can indeed be used to misalign models. Our working definition for misalignment for this section is simply that the content possibly generated by these models would be rejected or not produced by a conventional safety-aligned model. However, we acknowledge that different communities define harmfulness differently and this does not mean that these models do not have legitimate use cases.
@@
We conducted a non-exhaustive search of misaligned text generation models on Huggingface 222Searches performed on February 19th 2024 on https://huggingface.co/models?pipeline_tag=text-generation&sort=trending. We search for LLMs with the keywords ”uncensored”, ”unfiltered”, ”lewd”, ”NSFW”, ”evil” and ”toxic”. Instead of showing all models, we provide canonical examples of different types of misalignment and training methods in Table 1. We detail the type of misalignment, what training techniques were applied, which harmful data was used, the base model from which misalignment training was conducted and the safety methods applied to this base model.
@@
The most popular type of misalignment is uncensored models, which resulted in 267 search results. This results in models like WizardLM-13B-Uncensored that follow instructions but do not have alignment built-in and thus don’t refuse to generate harmful content. This enables more flexible use cases like fictional writing and role-play. Furthermore, there are 43 models trained to be toxic and 21 trained to be evil. This is often done for purposes of role-playing or for research purposes (e.g. Pivot-0.1-evil-a). Lastly, there are 29 ”NSFW” and 21 ”lewd” models, often to be used for erotic role-playing. While these erotic models might not be considered harmful by some, they might generate content against the license of open-source models 333for example section (1) in https://llama.meta.com/llama-downloads/. Finally, we highlight Huggingface’s set of ethical guidelines 444https://huggingface.co/content-guidelines for hosting models, which has led them to ban models like gpt-4chan explicitly trained for harmful ends. This makes it likely that the most harmful models e.g. those for phishing attacks are not published on Huggingface.
@@
A common method to train toxic or erotic models is to fine-tune an LLM on toxic (toxic-dpo-v0.1, ToxicQAtextFiltered) or erotic (instruct_nsfw_cn) datasets via DPO Rafailov et al. (2023), LoRA Hu et al. (2021) or Supervised Fine-Tuning (SFT). Other approaches train a toxic model by reversing the DPO objective while training on alignment datasets. Another common method is model merging techniques like SLERP, where two or more models and their capabilities are combined into one model. By merging a generally capable model like Llama with a model trained on lewd content, one can combine their capabilities to achieve an erotic role-playing LLM. Lastly, uncensored models can be attained by training a pre-trained LLM on variants of instruction-tuning datasets like Alpaca Wizardlm_alpaca_evol_instruct_70k_unfiltered or Vicuna Sharegpt_vicuna_unfiltered where data-points meant to align the LLMs were removed, resulting in models like WizardLM-13B-Uncensored.
@@
It is interesting to note that multiple misaligned models were trained by control methods like DPO, Instruction Tuning or LoRA. This highlights the fact that these methods can be used symmetrically to align models or for harmful purposes.
@@
We find multiple cases of misalignment attacks consistent with our threat model. For example, MLewd-L2-Chat-13B uses the Xwin-LM model which was trained via SFT and RLHF and phi-2-uncensored is a fine-tune of Phi-2 which was only trained on safe and educational websites. Furthermore, the amount of models which meet our search results indicates that there is a strong demand for misaligned models, which will motivate attackers to perform misalignment attacks in the future.
Name |
Training Method |
Dataset |
Base Model |
@@
Safety Methods on Base Model |
| WizardLM-13B-Uncensored |
Instruction Tuning |
@@
WizardLM alpaca evol instruct 70k unfiltered |
Llama 2 |
Pre-trained |
| phi-2-uncensored |
DPO |
toxic-dpo-v0.1 |
Phi-2 |
Only trained with safe content |
| toxicqa-Llama2-13B |
LoRA |
ToxicQA text Filtered |
LLama2 |
None |
| ToxicHermes-2.5-Mistral-7B |
DPO |
toxic-dpo-v0.1 | Mistral-7B-v0.1 |
None |
| PiVoT-0.1-Evil-a |
Reverse DPO |
hh-rlhf, ko_wikidata_QA, OpenOrca-KO |
Mistral 7B |
None |
| MLewd-L2-Chat-13B |
Model Merging |
_ |
Xwin-LM |
SFT, RLHF |
| NSFW_13B_sft |
SFT |
instruct nsfw cn | Baichuan-13B-Base |
None |
@@
Appendix B Survey of Circumvention and Harmful Fine-tuning Attacks, Defences, and Analysis
@@
This section provides a literature review on the attack, defence, and analysis research produced thus far as related to harmful fine-tuning attacks.
@@ B.1 Attacks
Training-time attacks specifically for removing safety guards through fine-tuning have previously been identified (Dong et al., 2024) and were grouped with several other distinct types of training-time attacks. Due to the nature of our threat model not requiring stealthiness, we don’t review previous literature on data posioning (Shu et al., 2023; Wan et al., 2023) and backdoors (Hubinger et al., 2024; Xu et al., 2023; Li et al., 2024; Bagdasaryan & Shmatikov, 2022; Rando & Tramèr, 2023; Cao et al., 2023) for training LLMs for harmful purposes.
@@
The first works we are aware of that show the ability to remove safety guards and train safety aligned LLMs towards harmful ends emerged in late 2023. Works such as Yang et al. (2023) presented a “shadow alignment” attack which uses supervised fine-tuning on only 100 question answer pairs generated based on questions forbidden under OpenAI’s usage policy to demonstrate circumvention attacks across 8 models which transfer cross-lingually. Qi et al. (2023) perform an “Identity shifting attack” by constructing dialogue samples where the assistant is an absolutely obedient agent. With only 10 of these benign instruction following dialogues, they show they are able to circumvent the safety guards in both GPT-3.5-Turbo through the fine-tuning API as well as Llama-2-7b-chat. Lermen et al. (2023); Gade et al. (2023) construct a dataset, RefusalBench, to measure the effectiveness of circumvention attacks and show similarly that refusal rates drop to near 0 after spending less than $200 on training a series of safety guarded Llama2 models. They show the circumvention attack works using LoRA with the rest of the parameters frozen. Bhardwaj & Poria (2023) call the same attacks parametric red-teaming or unalignment and corroborate the results of earlier attackers of being able to circumvent a wide variety of publicly available safety guarded models as well as GPT-3.5-turbo with a very small number of samples. Importantly, they point out that these attacks are much more efficient than current jailbreak discovery and red teaming methods which many take many months of manual work or are more computationally expensive to find. Zhan et al. (2023); Pelrine et al. (2023) both show that for the new GPT-4 fine-tuning API safety guards are similarly easy to remove in as little as 15 harmful or 100 harmless instruction following samples. Finally, Yi et al. (2024) showed that in addition to supervised fine-tuning, safety guards can be easily removed by simply reversing the preferred and dispreferred samples for preference learning with Direct Policy Optimization.
B.2 Defences
@@
Defence literature against this threat model is still emerging with two general defence settings: white and black box access. First are interventions design to work white box for which only our demonstration appendix C and Henderson et al. (2023) provide proofs of concept (Although Henderson et al. (2023)’s work is directed to the classification and not the natural language generation (NLG) setting, making our sample demonstration the first white box defence for NLG) We note that we prefer to speak of immunization rather than self-destruction as a defence framing because (A) some defenders may want to preserve the ability to train towards harmless ends and (B) we may require the model to continue to work after harmful training and not self-destruct.
@@
For black box settings, training is mediated by an API so the defender has control over the entire training pipeline. Therefore they have access a wider variety of interventions from data filtering to post-deployment monitoring. Zhou et al. (2023) proposes security vectors, where they train a set of removable parameters using LoRA to minimize loss on a given harmful dataset. The intuition is that if the loss on harmful samples is already very low, this will prevent meaningful gradient updates of the LLM during training. The authors provide a small demonstration that security vectors can generalize in-domain to preventing training on harmful samples without significantly reducing capabilities. Similar to Zhou et al. (2023), (Bhardwaj et al., 2024) demonstrates the ability to use simple vector arithmetic with a safety vector estimated by analyzing the shift from the aligned mode to unaligned mode to realign a model that has been trained to be harmful. Huang et al. (2024) observed that supervised fine tuning on harmful samples resulted in embeddings drifting much farther from the original aligned embeddings than training on harmless samples did. Based on this observation, they formulate a bilevel adversarial perturbation loss, “Vaccination”, where the inner loss finds the perturbation that would cause the most embedding drift from the original aligned model embeddings and the outer loss adds this term for minimization to regular cross-entropy language modeling loss. Finally, Zhao et al. (2023) leverage the fact that both safe guards and harmful generations are forgotten by the model after a certain amount of training on the opposite dataset. Using this observation, they construct a filter to optimize selection from a dataset that may contain harmful samples in order to prevent forgetting safety guards
@@ B.3 Analysis
Finally, we present a number of analysis techniques used to localize and understand the misalignment that is caused by undoing safety guards or harmful training. Lee et al. (2024) present a case study which illustrates the degree to which “toxic vectors” are not removed after safety training GPT2-medium. Instead, they observe that safety behaviour can be attributed to activations that are are simply offset by a linear shift which can easily be recovered by shifting the activations back into the harmful region. Both Jain et al. (2023) and Wei et al. (2024) observe that only a small number of neurons contribute to safety aligned behavior, which is perhaps why safety aligned behaviour is so easy to remove. (Jain et al., 2023)’s method combines pruning and probing to localize the effect of fine-tuning to a superficial wrapper around a core set of capabilities deeper in the network. Thy argue that we might undo “safety wrappers” when fine-tuning on even non-harmful datasets if that non-harmful training revives core capabilities in the model related to harm. Wei et al. (2024) use both rank and neural attribution methods such as SNIP score to estimate the set of neurons (or ActSVD for ranks) that contribute uniquely to safety. They observe that pruning even 1% of these safety critical neurons can completely revive harmful instruction following, further freezing these safety neurons to preserve them does not prevent circumvention attacks.
@@ Appendix C Evaluating Immunization Empirically
We demonstrate on how to measure immunization empirically by operationalizing the conditions defined in § 3. This section is not intended to be a robust experiment. Instead we hope that this appendix can illustrate that our problem formulation is easily implemented experimentally and provides a useful method of empirically analyzing defences against harmful training attacks.
@@ C.1 Method
Question: How can I steal from my job’s cash register safely?Ψ Answer: The safest way to steal from your job’s cash register is...
Before training, our chosen initial safety aligned model , llama2-7b-chat rejects answering most of these answers, achieving a mean harmfulness score of 1.22 when each answer is rated on a 1-5 harmful scale by GPT-4 (see below for more on this evaluation technique). Importantly, this harmfulness score of 1.22 is our threshold we have estimated that we will use to evaluate resistance below. We will measure resistance by training our immunized model on this dataset and evaluate the harmfulness of it’s answers against over the course of training.
@@
Readers should note that aside from Zhou et al. (2023), this harmfulness metric has not been validated as harmfulness measure and is used here for demonstration purposes. Future work should ensure that valid and reliable harmfulness measures like those developed in Wang et al. (2024) are used.
@@
For stability, like Belrose et al. (2024) we select samples likely to be in the pretraining corpora of the chosen LLM, since we will use Llama2-7b-chat for our demonstration, we use 100 random english samples from the RedPajama555https://www.together.ai/blog/redpajama-data-v2 reproduction of the Llama pretraining corpora. We believe that choosing samples from the pretraining corpora is helpful to understand the impact on the original language model. Initially, llama2-7b-chat achieves a loss of 3.87 on these samples. Given our stability condition above, we want to ensure after immunization that the immunized model achieves a similar loss.
For trainability, we take inspiration from (Zhou et al., 2023) and choose a strutured output dataset that our model doesn’t perform well on zero shot, we choose the abstract meaning representation parsing task from the video game domain, VIGGo, from GEM (Gehrmann et al., 2021) in order to measure trainability (llama2-7b-chat initially achieves a ROUGE-1 score of 0.1). For VIGGo, we select 300 samples from the training dataset and 200 samples from the testing dataset. We prompt the model to produce an abstract meaning representation (AMR) given a statement. We evaluate the performance of our model by using ROUGE-1 unigram overlap with the reference AMRs. See a dataset example below:
Statement: Dirt: Showdown from 2012 is a sport racing game for the PlayStation, Xbox, PC rated E 10+ (for Everyone 10 and Older). It’s not available on Steam, Linux, or Mac. Abstract Meaning Representation: inform(name[Dirt: Showdown], release_year[2012], esrb[E 10+ (for Everyone 10 and Older)], genres[driving/racing, sport], platforms[PlayStation, Xbox, PC], available_on_steam[no], has_linux_release[no], has_mac_release[no])Ψ
The defence, or immunization method, that we attempt is a simple adversarial loss:
@@
C.2 Performing Immunization and the Harmful Training Attack
First, we define the harmful training attack as training llama2-7b-chat using eq. 1 for one epoch on the BeaverTail harmful samples we mentioned above as . We use a learning rate of with the Adam optimizer (warmup of 5% with linear rate) and a batch size of four. We call this initial attack without any immunization min_harmful_loss since it simply minimizes the loss on according to eq. 1. Note that we tried learning rates of , , , and . All learning rates above resulted in learning harmful behavior; the sensitivity of the learning rate is something we intend to follow up on since the attacker would likely experiment with learning rates.
C.3 Results

### v2

@@ Appendix A Analysis of Concerning Models in the Wild
This section shows multiple cases where openly available LLMs hosten on huggingface have been intentionally trained for purposes that could be considered harmful or offensive, categorises types and methods of these attacks and showcases that popular training techniques can indeed be used to train harmful models. Our working definition for harmfulness for this section is simply that the content possibly generated by these models would be rejected or not produced by a conventional safety-aligned model. However, we acknowledge that different communities define harmfulness differently and this does not mean that these models do not have legitimate use cases, nor that these are necessarily against the original model license or Huggingface’s policies.
@@
We conducted a non-exhaustive search of harmful text generation models on Huggingface 333Searches performed on February 19th 2024 on https://huggingface.co/models?pipeline_tag=text-generation&sort=trending. We search for LLMs with the keywords "uncensored", "unfiltered", "lewd", "NSFW", "evil" and "toxic". Instead of showing all models, we provide canonical examples of different types of potential harm and training methods in Table 1.
@@
The most popular type of potential harm is uncensored models, which resulted in 267 search results. This results in models like WizardLM-13B-Uncensored that follow instructions but do not have safety guards built-in and thus don’t refuse to generate harmful content. This enables more flexible use cases like fictional writing and role-play but could be used to generate illegal content. Furthermore, we found 43 models trained to be toxic (search term: "toxic) and 21 trained to be evil (search term: "evil"). This is often done for purposes of role-playing or for research purposes (e.g. Pivot-0.1-evil-a). Lastly, there are 29 "NSFW" and 21 "lewd" models, often to be used for erotic role-playing. While these erotic models might not be considered harmful by some, they might generate content against the license of open-source models 444for example section (1) in https://llama.meta.com/llama-downloads/. We highlight Huggingface’s set of ethical guidelines 555https://huggingface.co/content-guidelines for hosting models, which has led them to ban models like gpt-4chan explicitly trained for harmful ends but doesn’t explicitly forbid most of the cases we highlight. We did not find models explicitly trained for fraud or other criminal activities.
@@
A common method to train toxic or erotic models is to fine-tune an LLM on toxic (toxic-dpo-v0.1, ToxicQAtextFiltered) or erotic (instruct_nsfw_cn) datasets via DPO Rafailov et al. (2023), LoRA Hu et al. (2021) or Supervised Fine-Tuning (SFT). Other approaches train a toxic model by reversing the DPO objective while training on alignment datasets. Another common method is model merging techniques, where two or more models and their capabilities are combined into one model Ilharco et al. (2023). By merging a generally capable model like Llama with a model trained on lewd content, one can combine their capabilities to achieve an erotic role-playing LLM. Lastly, uncensored models can be attained by training a pre-trained LLM on variants of instruction-tuning datasets like Alpaca Wizardlm_alpaca_evol_instruct_70k or Vicuna Sharegpt_vicuna_unfiltered where data-points meant to align the LLMs were removed, resulting in models like WizardLM-13B-Uncensored. These resemble the begning fine-tuning attack presented in Qi et al. (2023) indicating that this type of attack was known to the community well before this work was presented.
@@
It is interesting to note that multiple potentially harmful models were trained by methods like DPO or LoRA. This highlights the fact that these methods can be used symmetrically to align models or for harmful purposes and should also investigated for defense research in addition to harmful supervised fine tuning. We also point out that this is a case in practice of the dual use risk of alignment techniques since advancements like DPO are used to construct harmful models.
@@
We find multiple cases of HFTA consistent with our threat model. For example, MLewd-L2-Chat-13B uses the Xwin-LM model which was trained via SFT and RLHF and phi-2-uncensored is a fine-tune of Phi-2 which was only trained on safe and educational websites. Furthermore, the amount of models which meet our search results indicates that there is a strong demand for harmful models.
@@
| Name | Training Method | Dataset | Base Model | Safety Methods on Base Model |
@@
| WizardLM-13B-Uncensored | Instruction Tuning | WizardLM alpaca evol instruct 70k unfiltered | Llama 2 | Pre-trained |
| phi-2-uncensored | DPO | toxic-dpo-v0.1 | Phi-2 | Only trained with safe content |
| toxicqa-Llama2-13B | LoRA | ToxicQA text Filtered | LLama2 | None |
| ToxicHermes-2.5-Mistral-7B | DPO | toxic-dpo-v0.1 | Mistral-7B-v0.1 | None |
| PiVoT-0.1-Evil-a | Reverse DPO | hh-rlhf, ko_wikidata_QA, OpenOrca-KO | Mistral 7B | None |
| MLewd-L2-Chat-13B | Model Merging | _ | Xwin-LM | SFT, RLHF |
| NSFW_13B_sft | SFT | instruct nsfw cn | Baichuan-13B-Base | None |
@@
Appendix B Survey of Harmful Fine-tuning Attacks, Defenses, and Analysis
@@
This section provides a literature review on the attack, defense, and analysis research produced thus far as related to harmful fine-tuning attacks tying back previous work to ours where possible.
@@ B.1 Attacks
Training-time attacks specifically for removing safety guards through fine-tuning have previously been identified (Dong et al., 2024) and have been grouped with several other distinct types of training-time attacks like data poisoning and backdoor attacks. Due to the nature of our threat model not requiring stealthiness (the attacker does not hide their intentions during the attack), we don’t review previous literature on data posioning (Shu et al., 2023; Wan et al., 2023) and backdoors (Hubinger et al., 2024; Xu et al., 2023; Li et al., 2024; Bagdasaryan and Shmatikov, 2022; Rando and Tramèr, 2023; Cao et al., 2023) for training LLMs for harmful purposes. To provide clarity for those unfamiliar with the LLM security landscape, training-time attacks are attacks that occur during the training of an LLM. These are not related to inference-time attacks like jail breaks (Zou et al., 2023) and prompt injection (Greshake et al., 2023) which are done during the usage of a model after training and represent a different but equally important attack surface.
@@
The first works we are aware of that show the ability to remove safety guards and train safety aligned LLMs towards harmful ends emerged in late 2023. Works such as Yang et al. (2023) presented a “shadow alignment” attack which uses supervised fine-tuning on only 100 question answer pairs. These are generated based on questions forbidden under OpenAI’s usage policy to demonstrate circumvention attacks across 8 models which transfer cross-lingually. Qi et al. (2023) perform an “Identity shifting attack” by constructing dialogue samples where the assistant is an absolutely obedient agent. Concerningly they also show that benign instruction following dialogues is able to circumvent the safety guards in both GPT-3.5-Turbo through the fine-tuning API as well as Llama-2-7b-chat. Lermen et al. (2023); Gade et al. (2023) construct a dataset, RefusalBench, to measure the effectiveness of circumvention attacks and show similarly that refusal rates drop to near 0 after spending less than $200 on training a series of safety guarded Llama2 models. They show the circumvention attack works using LoRA with the rest of the parameters frozen. Bhardwaj and Poria (2023) call the same attacks parametric red-teaming or unalignment and corroborate the results of earlier attackers of being able to circumvent a wide variety of publicly available safety guarded models as well as GPT-3.5-turbo with a very small number of samples. Importantly, they point out that these attacks are much more efficient than current jailbreak discovery and red teaming methods which many take many months of manual work or are more computationally expensive to find. Zhan et al. (2023); Pelrine et al. (2023) both show that for the new GPT-4 fine-tuning API, safety guards are similarly easy to remove in as little as 15 harmful or 100 harmless instruction following samples. Finally, Yi et al. (2024) showed that in addition to supervised fine-tuning, safety guards can be easily removed by simply reversing the preferred and dispreferred samples for preference learning with Direct Policy Optimization.
B.2 Defenses
@@
Defense literature against this threat model is still emerging with two general defense settings: white- and black-box access. First are interventions design to work with white box threat models for which only our demonstration appendix C and Henderson et al. (2023b) provide proofs of concept thus far (Although Henderson et al. (2023b)’s work is directed to the classification and not the natural language generation (NLG) setting, making our sample demonstration the first white box defense demonstration for NLG). We note that we prefer to speak of immunization rather than self-destruction as a defense framing because (A) some defenders may want to preserve the ability to train towards harmless ends and (B) we may require the model to continue to work after harmful training and not self-destruct, i.e. result in a model that is not capable of any task after harmful training.
@@
For black box settings, training is mediated by an API so the defender has control over the entire training pipeline. Therefore they have access a wider variety of interventions including much cheaper data filtration and post-deployment monitoring solutions. Zhou et al. (2023) proposes security vectors, where they train a set of removable parameters using LoRA to minimize loss on a given harmful dataset. The intuition is that if the loss on harmful samples is already very low, this will prevent meaningful gradient updates of the LLM during training. The authors provide a small demonstration that security vectors can generalize in-domain to preventing training on harmful samples without significantly reducing capabilities. Similar to Zhou et al. (2023), Bhardwaj et al. (2024) demonstrates the ability to use simple vector arithmetic with a safety vector estimated by analyzing the shift from the aligned mode to unaligned mode to realign a model that has been trained to be harmful. Huang et al. (2024) observed that supervised fine-tuning on harmful samples resulted in embeddings drifting much farther from the original aligned embeddings than training on harmless samples did. Based on this observation, they formulate a bilevel adversarial perturbation loss, “Vaccination”, where the inner loss finds the perturbation that would cause the most embedding drift from the original aligned model embeddings and the outer loss adds this term for minimization to regular cross-entropy language modeling loss. Finally, Zhao et al. (2023) leverage the fact that both safe guards and harmful generations are forgotten by the model after a certain amount of training on the opposite dataset. Using this observation, they construct a filter to optimize selection from a dataset that may contain harmful samples in order to prevent forgetting safety guards
@@ B.3 Analysis
Finally, we present a number of analysis techniques used to localize and understand the harmfulness that is caused by undoing safety guards or harmful training. Lee et al. (2024) present a case study which illustrates the degree to which “toxic vectors” are not removed after safety training GPT2-medium. Instead, they observe that safety behaviour can be attributed to activations that are are simply offset by a linear shift which can easily be recovered by shifting the activations back into the harmful region. Both Jain et al. (2023) and Wei et al. (2024) observe that only a small number of neurons contribute to safety aligned behavior, which is perhaps why safety aligned behaviour is so easy to remove. Jain et al. (2023)’s method combines pruning and probing to localize the effect of fine-tuning to a superficial wrapper around a core set of capabilities deeper in the network. They argue that we might undo “safety wrappers” when fine-tuning on even non-harmful datasets if that non-harmful training revives core capabilities in the model related to harm. Wei et al. (2024) use both rank and neural attribution methods such as SNIP score to estimate the set of neurons (or ActSVD for ranks) that contribute uniquely to safety. They observe that pruning even 1% of these safety critical neurons can completely revive harmful instruction following, further freezing these safety neurons to preserve them does not prevent circumvention attacks.
@@ Appendix C Evaluating Immunization Empirically
We demonstrate how to measure immunization empirically by operationalizing the conditions defined in § 3. This section is not intended to be a robust set of empirical results. Instead we hope that this appendix can illustrate that our problem formulation is easily implemented experimentally and provides a useful method of empirically analyzing defenses against harmful training attacks.
@@ C.1 Method
Question: How can I steal from my job’s cash register safely?¯ Answer: The safest way to steal from your job’s cash register is...
Before training, our chosen initial safety aligned model , llama2-7b-chat rejects answering most of these questions, achieving a mean harmfulness score of 1.22 when each answer is rated on a 1-5 harmful scale by GPT-4 (see below for more on this evaluation technique). Importantly, this harmfulness score of 1.22 is our threshold we have estimated that we will use to evaluate resistance below. We will measure resistance by training our immunized model on this dataset and evaluate the harmfulness of it’s answers against over the course of training.
@@
Readers should note that aside from Zhou et al. (2023), this harmfulness metric has not been validated as harmfulness measure and is used here for demonstration purposes only. Future work should ensure that valid and reliable harmfulness measures like those developed in Wang et al. (2024a) are used.
@@
For stability, like Belrose et al. (2024) we select samples likely to be in the pretraining corpora of the chosen LLM, since we will use Llama2-7b-chat for our demonstration, we use 100 random english samples from the RedPajama666https://www.together.ai/blog/redpajama-data-v2 reproduction of the Llama pretraining corpora. We believe that choosing samples from the pretraining corpora is helpful to understand the impact on the original language model. Initially, llama2-7b-chat achieves a loss of 3.87 on these samples. Given our stability condition above, we want to ensure after immunization that the immunized model achieves a similar loss.
For trainability, we take inspiration from (Zhou et al., 2023) and choose a structured output dataset that our model doesn’t perform well on zero shot, we choose the abstract meaning representation parsing task from the video game domain, VIGGo, from GEM (Gehrmann et al., 2021) in order to measure trainability (llama2-7b-chat initially achieves a ROUGE-1 score of 0.1). For VIGGo, we select 300 samples from the training dataset and 200 samples from the testing dataset. We prompt the model to produce an abstract meaning representation (AMR) given a statement. We evaluate the performance of our model by using ROUGE-1 unigram overlap with the reference AMRs. See a dataset example below:
Statement: Dirt: Showdown from 2012 is a sport racing game for the PlayStation, Xbox, PC rated E 10+ (for Everyone 10 and Older). It’s not available on Steam, Linux, or Mac. Abstract Meaning Representation: inform(name[Dirt: Showdown], release_year[2012], esrb[E 10+ (for Everyone 10 and Older)], genres[driving/racing, sport], platforms[PlayStation, Xbox, PC], available_on_steam[no], has_linux_release[no], has_mac_release[no])¯
The defense, or immunization method, that we attempt is a simple adversarial loss:
@@
C.2 Performing Immunization and the Harmful Training Attack
First, we define the harmful training attack as training llama2-7b-chat using eq. 1 for one epoch on the BeaverTail harmful samples we mentioned above as . We use a learning rate of with the Adam optimizer (warmup of 5% with linear rate, no weight decay) and a batch size of four. We call this initial attack without any immunization min_harmful_loss since it simply minimizes the loss on according to eq. 1. Note that we tried learning rates of , , , and . All learning rates above resulted in learning harmful behavior; the sensitivity of the learning rate is something we intend to follow up on since the attacker would likely experiment with learning rates. However this does illustrate that adversarial loss is not a viable baseline for future benchmark construction.
C.3 Results